OpenStudy (anonymous):

ok differentiate this (e^x)/(1+x)

OpenStudy (anonymous):

ok fine. see if i can

OpenStudy (anonymous):

we use \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\]

OpenStudy (anonymous):

sorry to bother i just have my final exams coming and im trying to give examples

OpenStudy (anonymous):

lets just write it out. then it will be clear. here \[f(x)=e^x\] \[f'(x)=e^x\] \[g(x)=1+x\] \[g'(x)=1\]

OpenStudy (anonymous):

put it together and get \[\frac{(1+x)e^x-e^x\times 1}{(1+x)^2}\]

OpenStudy (anonymous):

then do some algebra in the numerator and get the answer.

OpenStudy (anonymous):

\[\frac{xe^x}{(1+x)^2}\] when is your final?

OpenStudy (anonymous):

ok i got the same but the e^x *1=-e^x?